3.2 g of oxygen (At. weight = 16) and 0.2 g of hydrogen (At. wt. = 1) are placed in a 1.12 litre flask at 0°C. The total pressure of the gas mixture will be


(a) 1 atm


(b) 4 atm


(c) 3 atm


(d) 2 atm


choose and explain

Grade:12

2 Answers

KUMAR PRASSU
11 years ago

 

There is a total of 0.2 moles of gas =>n=0.2

volume of container is 1.12L=>V=1.12L

temp. is 273 K=>T=273K

there fore by PV=nRT

pressure=4atm

there fore option 2 is correct

hope u like this 

PLS APPROVESmile

Vikas TU
11 years ago

You are dealing with diatomic oxygen ( molar mass 32g/mol ) and diatomic hydrogen ( 2g/mol )
Mol O2 in flask = 3.2/32 = 0.1 mol
Mol H2 in flask = 0.2/2 = 0.1 mol H2
Total moles = 0.2mol

Calculate pressure using gas equation:
PV = nRT
P * 1.12 = 0.2*0.082057*273
P = 4.480/1.12
P = 4.000 atm

Pressure = 4atm - choice b) is correct.

 

Plz Approve!